OpenStudy (anonymous):

add & simplify. (7/x^2-x-6) + 5/x^2-3

OpenStudy (anonymous):

((7)/(x^(2))-x-6)+(5)/(x^(2))-3 To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is x^(2). Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. (-x*(x^(2))/(x^(2))+(7)/(x^(2))-6*(x^(2))/(x^(2)))+(5)/(x^(2))-3 Complete the multiplication to produce a denominator of x^(2) in each expression. (-(x^(3))/(x^(2))+(7)/(x^(2))-(6x^(2))/(x^(2)))+(5)/(x^(2))-3 Combine the numerators of all expressions that have common denominators. ((-x^(3)+7-6x^(2))/(x^(2)))+(5)/(x^(2))-3 Reorder the polynomial -x^(3)+7-6x^(2) alphabetically from left to right, starting with the highest order term. ((-x^(3)-6x^(2)+7)/(x^(2)))+(5)/(x^(2))-3 Multiply the factor by the rest of the expression to remove the fraction from the denominator. To multiply by a factor in the denominator, multiply by 1 over the factor. (-x^(3)-6x^(2)+7)/(x^(2))+(5)/(x^(2))-3 To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is x^(2). Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. (-x^(3)-6x^(2)+7)/(x^(2))+(5)/(x^(2))-3*(x^(2))/(x^(2)) Complete the multiplication to produce a denominator of x^(2) in each expression. (-x^(3)-6x^(2)+7)/(x^(2))+(5)/(x^(2))-(3x^(2))/(x^(2)) Combine the numerators of all expressions that have common denominators. (-x^(3)-6x^(2)+7)/(x^(2))+(5-3x^(2))/(x^(2)) Reorder the polynomial 5-3x^(2) alphabetically from left to right, starting with the highest order term. (-x^(3)-6x^(2)+7)/(x^(2))+(-3x^(2)+5)/(x^(2)) The numerators of expressions that have equal denominators can be combined. In this case, ((-x^(3)-6x^(2)+7))/(x^(2)) and ((-3x^(2)+5))/(x^(2)) have the same denominator of x^(2), so the numerators can be combined. ((-x^(3)-6x^(2)+7)+(-3x^(2)+5))/(x^(2)) Simplify the numerator of the expression. (-x^(3)-6x^(2)+7-3x^(2)+5)/(x^(2)) Combine all similar terms in the polynomial -x^(3)-6x^(2)+7-3x^(2)+5. (-x^(3)-9x^(2)+12)/(x^(2))

OpenStudy (anonymous):

....? what is all this

OpenStudy (anonymous):

=(7/x^2-x-6) + 5/x^2-x-6 =12/x2-x-6 = ????

OpenStudy (anonymous):

is it 2/x^2-x-1